Zang Industries has hired the investment banking firm of Eric, Schwartz, & Mann (ESM) to help it go public. Zang and ESM agree that Zang's current value of equity is $62 million. Zang currently has 5 million shares outstanding and will issue 1.7 million new shares. ESM charges a 5% spread.

What is the correctly valued offer price? Do not round intermediate calculations. Round your answer to the nearest cent.

How much cash will Zang raise net of the spread (use the rounded offer price)? Write out your answer completely. For example, 5 million should be entered as 5,000,000. Round your answer to the nearest dollar.
Answer #1

a). Offer Price = Value of Equity / [(New Shares * Spread) + Old Shares]

= 62 / [(1.7 * 0.05) + 5]

= 62 / 5.085 = 12.19, or $12,190,000

b). Offer Price = Share Price * (1 - F)

= 12.19 * (1 - 0.05) = 11.58, or $11,580,000